Find new ammount.
Original Price- $14.57
Markup Percent- 105%

Answer by jim_thompson5910(35256)

105% of $14.57 = (105/100)*14.57 = 15.2985 = 15.30 (rounding to the nearest penny)

So the new price is 14.57+15.30 = $29.87